# Consistent eccentricities for gravitational wave astronomy: Resolving discrepancies between astrophysical simulations and waveform models

### v2

@@ Abstract
I Introduction
@@
Definitive measurement of orbital eccentricity in a compact binary coalescence (CBC) signal is one of the most exciting prospects of gravitational wave (GW) astrophysics. Despite $\mathcal{O}(100)$ binary black hole (BBH) signals being detected by the LIGO–Virgo–KAGRA (LVK) network LIGO Scientific Collaboration et al. (2015); Acernese et al. (2015); Aso et al. (2013); Akutsu et al. (2021); Abbott et al. (2018) to date Abbott et al. (2023a), the distinct pathways that led to the formation and merger of these BBH signals remains a mystery. Measurements of masses, spins, and redshift evolution have helped to identify broad constraints on the underlying physical processes that lead to BBH formation Abbott et al. (2023b), although measurement uncertainties paired with the inherent uncertainties embedded within formation channel modeling make it difficult to use these measurables to robustly pin down the formation pathway for a single BBH or the relative fraction of BBHs that results from one formation channel compared to another (see e.g. Zevin et al. (2021a); Cheng et al. (2023)).
@@
Measurable orbital eccentricity in stellar-mass BBH mergers, on the other hand, has been shown to be a robust indicator that a particular BBH signal resulted from a small subset of viable channels O’Leary et al. (2009); Kocsis and Levin (2012); Samsing et al. (2014); Antonini et al. (2017); Samsing and Ramirez-Ruiz (2017); Silsbee and Tremaine (2017); Rodriguez et al. (2018a); Rodriguez and Antonini (2018); Samsing (2018); Samsing et al. (2018a, b); Fragione and Kocsis (2019); Fragione and Bromberg (2019); Liu et al. (2019); Liu and Lai (2019); Zevin et al. (2019); Michaely and Perets (2020); Gondán and Kocsis (2021); Tagawa et al. (2021); Mapelli et al. (2021); Zevin et al. (2021b); Zhang et al. (2021); Samsing et al. (2022); Trani et al. (2022); Dall’Amico et al. (2023); Romero-Shaw et al. (2023a). Eccentricity is efficiently damped as a compact binary system inspirals and loses energy and angular momentum through GW emission Peters (1964). To retain any measurable semblance of orbital eccentricities at frequencies accessible to current ground-based GW detectors, a BBH system therefore must be able to form at a high eccentricity ($\gtrsim 0.9$) and merge on an extremely rapid timescale ($\lesssim 1\leavevmode\nobreak\ \text{yr}$, see Ref. Zevin et al. (2019)) or have some mechanism that continually pumps eccentricity into an inspiraling system (e.g., the von Zeipel-Lidov-Kozai mechanism, von Zeipel (1910); Lidov (1962); Kozai (1962)).
@@
The former of these possibilities is something that can only be attained through strong dynamical encounters within dense stellar environments, such as globular clusters, nuclear clusters, and the disks of active galactic nuclei. The synthesis of eccentric BBHs mergers in globular clusters is a robust and unavoidable result of simple physical processes such as two-body relaxation and (chaotic) gravitational encounters between three or more black holes within the cluster core (see e.g. Rodriguez et al. (2018a)). Since the relative fraction of systems that proceed through this eccentric subchannel is relatively robust to uncertainties in cluster environments, the detection and measurement of even a small number of eccentric BBH mergers can lead to stringent constraints on the relative fraction of merging BBH systems that result from dynamical environments entirely Zevin et al. (2021b).
@@
Although eccentricity is a powerful constraint on formation pathways in theory, in practice the measurement and interpretation of an eccentric GW signal is much more muddled. Current matched-filter CBC searches applied to GW data Abbott et al. (2023b) only use aligned-spin and quasi-circular templates Sachdev et al. (2019); Davies et al. (2020); Aubin et al. (2021), and thus a selection effect exists against detecting eccentric GW signals. The selection function for eccentric systems due to this method for searching the data has yet to be adequately characterized (see discussion in Ref. Zevin et al. (2021b)). Waveform-independent searches for eccentric mergers have also been carried out Abbott et al. (2019); Abac et al. (2023) with no significant triggers; this is unsurprising, since these searches are typically only sensitive to short duration (i.e. high mass) systems. A number of eccentric waveform approximants have been built (e.g. Tanay et al. (2016); Huerta et al. (2018); Nagar et al. (2021); Ramos-Buades et al. (2022); Liu et al. (2022)), although much more work needs to be done to make them accurate over a large range of eccentricities. In particular, simultaneously modeling eccentricity and spin precession has proved to be challenging (see Ref. Liu et al. (2023) for a recent attempt [^1] See also Ref. Klein (2021) that constructs an approximant that includes eccentricity and precession for binaries detectable with LISA), and these properties have degenerate effects on the GW waveform making it difficult to distinguish between precessing and eccentric hypotheses especially for massive signals Romero-Shaw et al. (2023b); Xu and Hamilton (2023); Romero-Shaw et al. (2020). Lastly, the definition of eccentricity differs between GW waveform models. Eccentricity evolves during the inspiral of a binary, and hence can only be defined with reference to an epoch in the orbit of that binary. However, the definition of this reference epoch and of eccentricity itself is inconsistent across different state-of-the-art eccentric waveform approximants. There have been attempts towards bringing all these definitions on an equal footing Knee et al. (2022); Ramos-Buades et al. (2022); Bonino et al. (2023); Shaikh et al. (2023), allowing consistent comparison of eccentricities derived from diverse waveform approximants.
@@
A robust and consistent definition for eccentricity that is shared across waveform models and astrophysical models is imperative to yield any astrophysical conclusions in the event of an eccentric BBH detection. In this paper, we show that the definition of eccentricity currently employed in astrophysical models is different from the definition used in waveform models. We also paint a clearer picture as to where these discrepancies in the definition of eccentricity lie, and present a framework for determining a robust definition of eccentricity that can be utilized for self-consistent comparisons between predictions from the astrophysical modeling community and the GW waveform community.
@@
Our paper is structured as follows. In Section II, we review current methods that are used for defining eccentricity, both in astrophysical simulations and waveform modeling. Section III presents a standardized definition of eccentricity and a scheme for extracting this definition of eccentricity from astrophysical simulations for a direct comparison to GW waveforms. We also show how quoted eccentricity values significantly differ when a self-consistent definition of eccentricity is not used, investigate the robustness of our improved definition to higher-order post-Newtonian (PN) effects and spin, and discuss how our self-consistent picture of eccentricity affects the inferred fraction of measurably eccentric sources from astrophysical models. Finally, we provide a summary and concluding remarks in Section IV. Example python scripts are available at https://github.com/adivijaykumar/standard-eccentricity-scripts, and data products are available on Zenodo Vijaykumar et al. (2024).
I.1 Definitions of eccentricity used in this work
@@
    $e_{\rm W03}$: Eccentricities calculated from the prescription outlined in Wen (2003). This prescription has been commonly used to extract eccentricities from astrophysical simulations of binary formation (e.g. Samsing and Ramirez-Ruiz (2017); Samsing (2018); Rodriguez et al. (2018a); Zevin et al. (2019, 2021b); Kritos et al. (2022); Chattopadhyay et al. (2023)). As we shall see in Section II.2, this prescription uses the GW peak frequency as the reference frequency.
@@
    $e_{t}^{k{\rm PN}}$: The eccentricity computed by evolving PN equations accurate to the $k$-th order (see e.g. Tanay et al. (2016); Klein et al. (2018)). This uses 22 mode GW frequency as the reference frequency.
@@
    $e_{\rm W03}^{f_{22}}$: Eccentricities calculated using the W03 prescription, but extracted at fixed 22 mode frequency.
@@
    $e_{\rm gw}^{\rm wf}$: Eccentricity definition proposed in Shaikh et al. (2023), computed from a gravitational waveform using the gw_eccentricity package (see also Ref. Ramos-Buades et al. (2022)).
@@
II Background
II.1 Peak harmonic of a GW signal in presence of eccentricity
Figure 1: Numerical solution of Eq. (1) for the peak harmonic plotted along with the analytical fits from Ref. Wen (2003) and Ref. Hamers (2021). We see that the peak harmonic increases with eccentricity, and reaches values $>100$ for $e>0.9$. We also see that the fit in Ref. Hamers (2021) fits the numerical solution better, especially at lower values of eccentricity.
@@ II.1 Peak harmonic of a GW signal in presence of eccentricity
Note that the harmonic index $n$ is different from the $(\ell,m)$ basis used to parametrize GW polarizations. At leading order, power emitted by the $n$-th harmonic is given by Peters and Mathews (1963),
@@
where the function $g(n,e)$ quantifies the enhancement in the power emitted in the $n^{\mathrm{th}}$ harmonic compared to a circular orbit (where there is no emission at $n>2$), and is given by
@@
Here, $J_{k}(x)$ refers to Bessel functions of the first kind. For a given eccentricity, the peak power is emitted at a harmonic where $g(n,e)$ has the maximum value. For ease of calculation, Ref. Wen (2003) provided a fit to the peak harmonic as a function of eccentricity,
@@
Ref. Hamers (2021) noted that the above fit does not work well at low eccentricities, and provided an alternative fit to the peak harmonic,
@@
II.2 Eccentricity extraction prescription of Ref. Wen (2003)
Since the value of eccentricity changes during the course of the inspiral, the value of eccentricity is often quoted at a specific reference frequency (or reference time). Ref. Wen (2003) (henceforth referred to as W03) uses the peak harmonic’s frequency as a reference for defining eccentricity, and this definition has been employed by many other studies (e.g. Samsing and Ramirez-Ruiz (2017); Samsing (2018); Rodriguez et al. (2018a); Zevin et al. (2019, 2021b); Kritos et al. (2022); Chattopadhyay et al. (2023)). Given semimajor axis $a$ and total mass $M$, the average orbital frequency $f_{\mathrm{orb}}$ is
@@
In astrophysical simulations, we are generally interested in obtaining eccentricity at a specified reference frequency, given an initial value of semi-major axis $a_{0}$ and eccentricity $e_{0}$. These initial values can be thought of as arising either from some fiducial distribution of eccentricities and separations, or from stopping conditions in dynamical simulations. W03 uses the leading (Newtonian) order evolution Peters (1964) of the semi-major axis and eccentricity to relate $(a_{0},e_{0})$ to $(a,e)$ at a different time,
@@
II.3 Possible issues with the W03 prescription
@@
    The reference frequency is taken to be the peak frequency of the orbit. This is inconsistent with recommendations in the literature attempting to standardize the definition of eccentricity in waveforms Ramos-Buades et al. (2022); Shaikh et al. (2023). Specfically, these works motivate using the orbit-averaged 22 (i.e. $\ell=2$, $m=2$) mode frequency as a consistent reference point, on account of its smooth and monotonic evolution over the course of the inspiral.
@@
    Since all equations are based on Ref. Peters (1964), they are restricted to Newtonian (i.e. leading) order. This is not a good approximation of the evolution of the orbit, especially for systems close to merger. Specifically, the ordinary differential equations describing the evolution of eccentricities and the 22 mode frequencies have been calculated to several PN orders beyond the Newtonian order, also with the inclusion of spins. Furthermore, this equation breaks down at very high eccentricity.
In the rest of our paper, we shall formulate a procedure that can be applied to any set of scattering experiments to extract eccentricities, and conduct sanity checks of our procedure. In order to accomplish this, we shall solve the problems listed above [^2] Beyond the drawbacks mentioned here, we also note that, on account of being orbit-averaged, the Peters equations break down when the orbital time scale $\tau_{\rm orb}$ is much greater than the radiation reaction time scale $\tau_{\rm rr}$ (e.g. at large separations and high eccentricity; see for instance the discussion in Fumagalli and Gerosa (2023) and references therein.). However, since end states of binaries from typical simulations follow $\tau_{\rm rr}\gg\tau_{\rm orb}$, we do not attempt to account for this issue. .
III A consistent prescription for calculating eccentricity
As shown in Figure 1, the peak harmonic of an eccentric inspiral takes higher values for higher eccentricities. For low values of eccentricity ($e\lesssim 0.3$), the peak harmonic is $n_{\rm peak}=2$, and hence the 22-mode frequency is the same as frequency of the peak harmonic. However, the difference between the peak harmonic frequency and the 22-mode frequency increases very quickly, with $f_{\rm peak}=10f_{\rm orb}=5f_{22}$ at $e\approx 0.7$. This discrepancy means that the eccentricities derived from astrophysical simulations of different binary formation channels is defined at a different reference frequency as compared to the definition of reference frequency built into waveforms as per recommendations of Ref. Shaikh et al. (2023). As is also evident, this discrepancy is worse for higher eccentricities, which are arguably more interesting from an astrophysical standpoint.
To remedy this discrepancy, the definitions of eccentricity from the simulations and waveforms need to be placed on the same footing. Hence, instead of defining eccentricity at the frequency of the peak harmonic, we define eccentricity at the orbit-averaged 22 mode frequency. Let $M=m_{1}+m_{2}$ be the total mass of the binary and $\eta=(m_{1}m_{2})/M^{2}$ be the symmetric mass ratio, where $m_{1}$ and $m_{2}$ are the component binary masses. Given an initial value of semimajor axis $a_{0}$ and eccentricity $e_{0}$ for a particular binary from a simulation, we evolve differential equations in $e$ and $v$ given by PN formulae,

### v3

@@ Abstract
software: numpy (Harris et al., 2020), scipy (Virtanen et al., 2020), matplotlib (Hunter, 2007), astropy (Astropy Collaboration et al., 2013, 2018, 2022), jupyter (Kluyver et al., 2016), pandas (McKinney, 2010), seaborn (Waskom, 2021), and lalsuite (LIGO Scientific Collaboration et al., 2018).
1 Introduction
@@
Definitive measurement of orbital eccentricity in a compact binary coalescence (CBC) signal is one of the most exciting prospects of gravitational wave (GW) astrophysics. Despite $\mathcal{O}(100)$ binary black hole (BBH) signals being detected by the LIGO–Virgo–KAGRA (LVK) network (LIGO Scientific Collaboration et al., 2015; Acernese et al., 2015; Aso et al., 2013; Akutsu et al., 2021; Abbott et al., 2018) to date (Abbott et al., 2023a), the distinct pathways that led to the formation and merger of these BBH signals remains a mystery. Measurements of masses, spins, and redshift evolution have helped to identify broad constraints on the underlying physical processes that lead to BBH formation (Abbott et al., 2023b), although measurement uncertainties paired with the inherent uncertainties embedded within formation channel modeling make it difficult to use these measurables to robustly pin down the formation pathway for a single BBH or the relative fraction of BBHs that results from one formation channel compared to another (see e.g. Zevin et al. 2021b; Cheng et al. 2023).
@@
Measurable orbital eccentricity in stellar-mass BBH mergers, on the other hand, has been shown to be a robust indicator that a particular BBH signal resulted from a small subset of viable channels (O’Leary et al., 2009; Kocsis & Levin, 2012; Samsing et al., 2014; Antonini et al., 2017; Samsing & Ramirez-Ruiz, 2017; Silsbee & Tremaine, 2017; Rodriguez et al., 2018a; Rodriguez & Antonini, 2018; Samsing, 2018; Samsing et al., 2018a, b; Fragione & Kocsis, 2019; Fragione & Bromberg, 2019; Liu et al., 2019; Liu & Lai, 2019; Zevin et al., 2019; Michaely & Perets, 2020; Gondán & Kocsis, 2021; Tagawa et al., 2021; Mapelli et al., 2021; Zevin et al., 2021a; Zhang et al., 2021; Samsing et al., 2022; Trani et al., 2022; Dall’Amico et al., 2023; Romero-Shaw et al., 2023a; Trani et al., 2023). Eccentricity is efficiently damped as a compact binary system inspirals and loses energy and angular momentum through GW emission (Peters, 1964). To retain any measurable semblance of orbital eccentricities at frequencies accessible to current ground-based GW detectors, a BBH system therefore must be able to form at a high eccentricity ($\gtrsim 0.9$) and merge on an extremely rapid timescale ($\lesssim 1~{}\text{yr}$, see Zevin et al. 2019) or have some mechanism that continually pumps eccentricity into an inspiraling system (e.g., the von Zeipel-Lidov-Kozai mechanism von Zeipel 1910; Lidov 1962; Kozai 1962; Naoz 2016).
@@
The former of these possibilities is something that can only be attained through strong dynamical encounters within dense stellar environments, such as globular clusters, nuclear clusters, and the disks of active galactic nuclei. The synthesis of eccentric BBHs mergers in globular clusters is a robust and unavoidable result of simple physical processes such as two-body relaxation and (chaotic) gravitational encounters between three or more black holes within the cluster core (see e.g. Rodriguez et al. 2018a). Since the relative fraction of systems that proceed through this eccentric subchannel is relatively robust to uncertainties in cluster environments, the detection and measurement of even a small number of eccentric BBH mergers can lead to stringent constraints on the relative fraction of merging BBH systems that result from dynamical environments entirely (Zevin et al., 2021a).
@@
Although eccentricity is a powerful constraint on formation pathways in theory, in practice the measurement and interpretation of an eccentric GW signal is much more muddled. Current matched-filter CBC searches applied to GW data (Abbott et al., 2023b) only use aligned-spin and quasi-circular templates (Sachdev et al., 2019; Davies et al., 2020; Aubin et al., 2021), and thus a selection effect exists against detecting eccentric GW signals. The selection function for eccentric systems due to this method for searching the data has yet to be adequately characterized (see discussion in Zevin et al. 2021a). Waveform-independent searches for eccentric mergers have also been carried out (Abbott et al., 2019; Abac et al., 2023) with no significant triggers; this is unsurprising, since these searches are typically only sensitive to short duration (i.e. high mass) systems. A number of eccentric waveform approximants have been built (Tanay et al., 2016; Huerta et al., 2018; Nagar et al., 2021; Ramos-Buades et al., 2022; Liu et al., 2022), although much more work needs to be done to make them accurate over a large range of eccentricities. In particular, simultaneously modeling eccentricity and spin precession has proved to be challenging (see Liu et al. 2023 for a recent attempt [^1] See also Klein (2021) that constructs an approximant that includes eccentricity and precession for binaries detectable with LISA), and these properties have degenerate effects on the GW waveform making it difficult to distinguish between precessing and eccentric hypotheses especially for massive signals (Romero-Shaw et al., 2023b; Xu & Hamilton, 2023; Romero-Shaw et al., 2020). Lastly, the definition of eccentricity differs between GW waveform models. Eccentricity evolves during the inspiral of a binary, and hence can only be defined with reference to an epoch in the orbit of that binary. However, the definition of this reference epoch and of eccentricity itself is inconsistent across different state-of-the-art eccentric waveform approximants, in part because eccentricity is a gauge-dependent quantity within general relativity. There have been attempts towards bringing all these definitions on an equal footing (Knee et al., 2022; Ramos-Buades et al., 2022; Bonino et al., 2023; Shaikh et al., 2023), allowing consistent comparison of eccentricities derived from diverse waveform approximants.
@@
A robust and consistent definition for eccentricity that is shared across waveform models and astrophysical models is imperative to yield any astrophysical conclusions in the event of an eccentric BBH detection. In this paper, we show that the definition of eccentricity currently employed in astrophysical models is different from the definition used in waveform models. This discrepancy between these definitions is relevant for eccentricity $e\gtrsim 0.2$ with a factor of $\sim 2$ discrepancy at high $e$. We also paint a clearer picture as to where these discrepancies in the definition of eccentricity lie, and present a framework for determining a robust definition of eccentricity that can be utilized for self-consistent comparisons between predictions from the astrophysical modeling community and the GW waveform community.
@@
Our paper is structured as follows. In Section 2, we review current methods that are used for defining eccentricity, both in astrophysical simulations and waveform modeling. Section 3 presents a standardized definition of eccentricity and a scheme for extracting this definition of eccentricity from astrophysical simulations for a direct comparison to GW waveforms. We also show how quoted eccentricity values significantly differ when a self-consistent definition of eccentricity is not used, investigate the robustness of our improved definition to higher-order post-Newtonian (PN) effects and spin, and discuss how our self-consistent picture of eccentricity affects the inferred fraction of measurably eccentric sources from astrophysical models. Finally, we provide a summary and concluding remarks in Section 4. Example python scripts (Vijaykumar et al., 2024a) and data products (Vijaykumar et al., 2024b) are available on Zenodo.
1.1 Definitions of eccentricity used in this work
@@
    $e_{\rm W03}$: Eccentricities calculated from the prescription outlined in Wen (2003). This prescription has been commonly used to extract eccentricities from astrophysical simulations of binary formation (e.g. Samsing & Ramirez-Ruiz 2017; Samsing 2018; Rodriguez et al. 2018a; Zevin et al. 2019, 2021a; Kritos et al. 2022; Chattopadhyay et al. 2023). As we shall see in Section 2.2, this prescription uses the GW peak frequency as the reference frequency.
@@
    $e_{t}^{k{\rm PN}}$: The eccentricity computed by evolving PN equations accurate to the $k$-th order (see e.g. Tanay et al. 2016; Klein et al. 2018). This uses the dominant, quadrupolar mode ($\ell=m=2$) GW frequency as the reference frequency.
@@
    $e_{\rm W03}^{f_{22}}$: Eccentricities calculated using the W03 prescription, but extracted at fixed 22-mode frequency.
@@
    $e_{\rm gw}^{\rm wf}$: Eccentricity definition proposed in Shaikh et al. (2023), computed from a gravitational waveform using the gw_eccentricity package (see also Ramos-Buades et al. 2022).
@@
2 Background
2.1 Peak harmonic of a GW signal in presence of eccentricity
Figure 1: Numerical solution of Eq. (1) for the peak harmonic plotted along with the analytical fits from Wen (2003) and Hamers (2021). We see that the peak harmonic increases with eccentricity, and reaches values $>100$ for $e>0.9$. We also see that the fit in Hamers (2021) fits the numerical solution better, especially at lower values of eccentricity.
@@ 2.1 Peak harmonic of a GW signal in presence of eccentricity
Note that the harmonic index $n$ is different from the $(\ell,m)$ basis used to parametrize GW polarizations. At leading order, power emitted by the $n$-th harmonic is given by (Peters & Mathews, 1963),
@@
where the function $g(n,e)$ quantifies the enhancement in the power emitted in the $n^{\mathrm{th}}$ harmonic compared to a circular orbit (where there is no emission at $n\neq 2$), and is given by
@@
Here, $J_{k}(x)$ refers to Bessel functions of the first kind. For a given eccentricity, the peak power is emitted at a harmonic where $g(n,e)$ has the maximum value. For ease of calculation, Wen (2003) provided a fit to the peak harmonic as a function of eccentricity,
@@
Hamers (2021) noted that the above fit does not work well at low eccentricities, and provided an alternative fit to the peak harmonic,
@@
2.2 Eccentricity extraction prescription of Wen (2003)
Since the value of eccentricity changes during the course of the inspiral, the value of eccentricity is often quoted at a specific reference frequency (or reference time). Wen (2003) (henceforth referred to as W03) uses the peak harmonic’s frequency as a reference for defining eccentricity, and this definition has been employed by many other studies (e.g. Samsing & Ramirez-Ruiz 2017; Samsing 2018; Rodriguez et al. 2018a; Zevin et al. 2019, 2021a; Kritos et al. 2022; Chattopadhyay et al. 2023). Given semimajor axis $a$ and total mass $M$, the average orbital frequency $f_{\mathrm{orb}}$ is
@@ 2.2 Eccentricity extraction prescription of Wen (2003)
In astrophysical simulations, we are generally interested in obtaining eccentricity at a specified reference frequency, given an initial value of semi-major axis $a_{0}$ and eccentricity $e_{0}$. These initial values can be thought of as arising either from some fiducial distribution of eccentricities and separations, or from stopping conditions in dynamical simulations. W03 uses the leading (Newtonian) order evolution (Peters, 1964) of the semi-major axis and eccentricity to relate $(a_{0},e_{0})$ to $(a,e)$ at a different time, {widetext}
@@
2.3 Possible issues with the W03 prescription
@@
    The reference frequency is taken to be the peak frequency of the orbit. This is inconsistent with recommendations in the literature attempting to standardize the definition of eccentricity in waveforms (Ramos-Buades et al., 2022; Shaikh et al., 2023). Specfically, these works motivate using the orbit-averaged 22 (i.e. $\ell=2$, $m=2$) mode frequency as a consistent reference point, on account of its smooth and monotonic evolution over the course of the inspiral.
@@
    Since all equations are based on Peters (1964), they are restricted to Newtonian (i.e. leading) order. This is not a good approximation of the evolution of the orbit, especially for systems close to merger. Specifically, the ordinary differential equations describing the evolution of eccentricities and the 22 mode frequencies have been calculated to several PN orders beyond the Newtonian order, also with the inclusion of spins. Furthermore, this equation breaks down at very high eccentricity.
In the rest of our paper, we shall formulate a procedure that can be applied to any set of scattering experiments to extract eccentricities, and conduct sanity checks of our procedure. In order to accomplish this, we shall solve the problems listed above [^2] Beyond the drawbacks mentioned here, we also note that, on account of being orbit-averaged, the Peters equations break down when the orbital time scale $\tau_{\rm orb}$ is much greater than the radiation reaction time scale $\tau_{\rm rr}$ (e.g. at large separations and high eccentricity; see for instance the discussion in Fumagalli & Gerosa 2023 and references therein.). However, since end states of binaries from typical simulations follow $\tau_{\rm rr}\gg\tau_{\rm orb}$, we do not attempt to account for this issue. .
3 A consistent prescription for calculating eccentricity
As shown in Figure 1, the peak harmonic of an eccentric inspiral takes higher values for higher eccentricities. For low values of eccentricity ($e\lesssim 0.3$), the peak harmonic is $n_{\rm peak}=2$, and hence the 22-mode frequency is the same as frequency of the peak harmonic. However, the difference between the peak harmonic frequency and the 22-mode frequency increases very quickly, with $f_{\rm peak}=10f_{\rm orb}=5f_{22}$ at $e\approx 0.7$. This discrepancy means that the eccentricities derived from astrophysical simulations of different binary formation channels is defined at a different reference frequency as compared to the definition of reference frequency built into waveforms as per recommendations of Shaikh et al. (2023). As is also evident, this discrepancy is worse for higher eccentricities, which are arguably more interesting from an astrophysical standpoint.
To remedy this discrepancy, the definitions of eccentricity from the simulations and waveforms need to be placed on the same footing. Hence, instead of defining eccentricity at the frequency of the peak harmonic, we define eccentricity at the orbit-averaged 22 mode frequency. Let $M=m_{1}+m_{2}$ be the total mass of the binary and $\eta=(m_{1}m_{2})/M^{2}$ be the symmetric mass ratio, where $m_{1}$ and $m_{2}$ are the component binary masses. Given an initial value of semimajor axis $a_{0}$ and eccentricity $e_{0}$ for a particular binary from a simulation, we evolve differential equations in $e$ and $v$ given by PN formulae, {widetext}